Solve the initial-value problem : (dx)/(dt) +(tant)x=cos^2(t), x(0)=-1
1
Expert's answer
2021-07-09T14:18:33-0400

Integrating factor: "\mu(t)=\sec t"


"\sec t\dfrac{dx}{dt}+\sec t(\tan t)x=\cos t"

"\dfrac{d}{dt}(x\sec t)=\cos t"

"x\sec t=\int\cos t dt"

"x\sec t=\sin t+C"

"x(0)=-1"


"-1\cdot\sec(0)=\sin (0)+C"


"C=-1"

"x\sec t=\sin t-1"

Or


"x=\sin t\cos t-\cos t"